a hollow metal sphere of mass m has an instantaneous upward acceleration a when released from rest fully submerged in a liquid of density d1.if the density of the metal is d2 find the volume of cavity in the sphere 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
The eqn. of force Balane is:
B – mg = ma
B= mg + ma
(d1*V*g) = m(g+a)
V = m(g+a)/d1*g would be the Volume of the cavity.